# On Picard’s Theorem Via Nevanlinna Theory

### v2

@@ Abstract.
We study a long-standing Picard’s problem for non-compact complete Kähler manifolds with non-negative Ricci curvature. When the manifold carries a positive Green function, a positive answer is given to the Picard’s problem, i.e., it is showed that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values.
@@ 1.1. Motivations and Techniques
It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values. All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold. Due to S. T. Yau [44], each holomorphic function on a non-compact complete Kähler manifold (Hermitian manifold, more general) with non-negative Ricci curvature has Liouville property. So, a natural problem is that does Picard’s little theorem still hold on such a manifold? It is called the Picard’s problem. To our knowledge, this is a long-standing problem. In what follows, we shall give an accurate statement.
@@
The first result in this direction could be traced back to 1970. S. Kobayashi [20] obtained some Picard-type theorems for holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity. As a special consequence, he showed that
@@
It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation. The first result without any restrictions for is due to S. T. Yau [44], who considered the Liouville’s problem based on a method of gradient estimates at the same time, which is viewed as a weaker version of the Picard’s problem. He showed that
@@
By S. T. Yau [44], Liouville’s theorem still holds for all harmonic functions on (see Cheng-Yau [9] also). The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem. Until 2010, A. Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate. In the following, we introduce his basic ideas. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with Fubini-Study metric Let be a meromorphic function. Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
@@
In the present paper, we wish to settle the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature. As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R. Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925. During this period, rich results have been achieved, we may refer the reader to L. V. Ahlfors [4], H. Cartan [6], Carlson-Griffths-King [7, 16], J. Noguchi [23, 24], E. I. Nochka [25], M. Ru [28, 29], B. Shiffman [30], B. Shabat [31], F. Sakai [32, 33], W. Stoll [37, 38], P. Vojta [36], H. Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and their references therein. To our knowledge, almost all theorems from Nevanlinna theory are based only on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we know very little if the domain manifolds are not non-positively curved, since we know little about how to estimate Green functions (satisfying Dirichelet boundary condition) for bounded domains under a Ricci curvature condition.
In 2023, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature. As a consequence, the author gave a positive answer to the Picard’s problem if is of maximal volume growth saying that
@@
The original technique in [14] is to construct a family of relatively compact domains exhausting through the global Green functions for based on an asymptotic estimate of minimal positive global Green function obtained by Colding-Minicozzi [8]. With an optimal estimate for local Green function for in terms of integral forms, the author (see [14]) established a Second Main Theorem of meromorphic mappings on Let us introduce the main contributions or key techniques in [14]. Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) obtained the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of the asymptotic estimate of the author defined the domain:
@@
According to [14], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma can be established. The situation where is trivial for that is conformally equivalent to under two assumptions of non-negative Ricci curvature and maximal volume growth (see [21]). By employing the standard arguments, the author established a Second Main Theorem of meromorphic mappings from into a complex projective manifold, which gives a Picard’s theorem:
@@
In this paper, we will give a positive answer under a much weaker volume growth condition:
@@
According to the estimate for obtained by Li-Yau [22], there exist constants such that
@@ 1.2. Main results
Let be a complex projective manifold with over which we can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper. Assume that carries a positive global Green function, i.e., it satisfies the volume growth condition:
Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined as
The first main result is the following Second Main Theorem.
@@
Let be a differentiably non-degenerate meromorphic mapping. Let be a reduced divisor of simple normal crossing type. Then for any there exists a subset of finite Lebesgue measure such that
@@
Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values. Actually, this theorem gives a quantitative solution to the Picard’s problem.
@@ 2.1. Volume Comparison Theorem
A space form is defined as a complete Riemannian manifold with constant sectional curvature. Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Denote by the Riemannian volume of a geodesic ball centered at with radius
Bishop-Gromov (see, e.g., [5]) gave an upper bound of :
@@
Calabi-Yau (see, e.g., [35]) gave a lower bound of :
@@
Assume that is non-compact. If then has an infinite volume. More precisely, there exists a constant such that
@@ 2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of for is the minimal positive fundamental solution of the following heat equation
@@ 3.1. Construction of
Assume that is satisfied the condition of volume growth:
It implies that is non-parabolic and there exists a unique minimal positive global Green function for satisfying
@@
It is evident that is relatively compact for any and the sequence exhausts if According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
@@ 3.2. Nevanlinna’s Functions
where is the volume element of
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential By the Kählerity of we have
@@ 3.3. First Main Theorem
To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) which states that
@@ 4. Gradient Estimates of Green Functions
Let be a non-compact complete Kähler manifold of complex dimension satisfying
@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-compact complete Kähler manifold of complex dimension satisfying
@@ 5.1. Calculus Lemma
Now, we apply Borel’s lemma to the left hand side of the above equality for twice: one is to and the other is to we conclude that for any there exists a subset of finite Lebesgue measure such that
@@ 5.2. Logarithmic Derivative Lemma
Assume that Then for any there exists a subset of finite Lebesgue measure such that
@@
The concavity of implies that
By this with Corollary 5.3, we show that for any there exists a subset of finite Lebesgue measure such that
holds for all outside Disturbing such that is absorbed. The proof is completed. ∎
@@
Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
@@ 6. Second Main Theorem and Defect Relation
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature satisfying
@@ 6.1. Second Main Theorem
in the sense of currents, where is the reduced divisor of and is the Ricci form of Therefore, it yields that
Let us establish a Second Main Theorem as follows:
@@
Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@ 6.2. Defect Relation
Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Denote by the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
@@
Corollary 6.8 (Picard Theorem).
@@ 6.3. Picard’s Theorem
Let be distinct values in Let be a nonconstant meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that

### v5

@@ Abstract.
We consider the Picard’s problem via Nevanlinna theory for complete Kähler manifolds with non-negative Ricci curvature. When the manifold is non-parabolic, which means that it carries a positive global Green function, we give a positive answer to the Picard’s problem. That is, we confirm that every meromorphic function on such a (non-compact) manifold reduces to a constant if it omits three distinct values.
@@ 1.1. Motivations and Techniques
It is well-known that the famous Picard’s little theorem asserts that every meromorphic function on the complex Euclidean space reduces to a constant, if it omits three distinct values. All the time, many authors have been trying to generalize this theorem to a more general non-compact complex manifold. Due to S. T. Yau [44], each holomorphic function on a non-compact complete Kähler manifold (or Hermitian manifolds in general) with non-negative Ricci curvature has Liouville property. So, a natural problem is that does Picard’s little theorem still hold on such a manifold? We call it the Picard’s problem. To our knowledge, it is still an unsolved problem. In what follows, we would like to give a precise statement.
@@
Remark. Let be the universal covering, where is the unit disc in It is not very hard to confirm that the Picard’s problem can be given a positive answer using Yau’s theorem (see Theorem C below), if any holomorphic mapping lifts to a holomorphic mapping via i.e., Using the criterion for the existence of the liftings, can be lifted to if and only if for some such that Since is simply-connected, then is trivial. This implies that such a lifting exists if and only if is trivial. However, is often not a trivial group, so one cannot lift to in general.
The first result in this direction could be traced back to 1970, S. Kobayashi [20] proved some Picard-type theorem for a holomorphic mappings between complex manifolds from the viewpoint of complex hyperbolicity. As a special consequence, he showed that
@@
It does not seem to be a quite satisfactory answer to the Picard’s problem, for some restrictions such as local flatness and bounded dilatation. The first result without any restrictions for is due to S. T. Yau [44], who considered the Liouville’s problem based on the gradient estimation method at the same time, which is viewed as a weaker version of the Picard’s problem. He showed that
@@
By S. T. Yau [44], Liouville’s theorem works for all the harmonic functions on (see Cheng-Yau [9] also). The gradient estimation method shows great power in many problems from geometric analysis (see, e.g., [35]), but it seems to be unable to give a solution to the Picard’s problem. Until 2010, A. Atsuji [2] developed a Nevanlinna-type theory based on heat diffusions, and he gave a positive answer to the Picard’s problem for as a meromorphic function on is of a slow growth rate. In the following, we introduce his basic ideas. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with Fubini-Study metric Let be a meromorphic function. Using the Kählerness of the Hilbert-Schmidt norm of differential (with respect to metrics ) can be written as the form
@@
In the present paper, we shall consider the Picard’s problem by developing Nevanlinna theory (see, e.g., [24, 26, 29]) to complete Kähler manifolds with non-negative Ricci curvature. As we know, Nevanlinna theory studies value distribution of meromorphic mappings between complex spaces, which has been developed for a long time since R. Nevanlinna founded two fundamental theorems for meromorphic functions on the complex plane in 1925. During this period, rich results have been achieved, we may refer the reader to L. V. Ahlfors [4], H. Cartan [6], Carlson-Griffths-King [7, 16], J. Noguchi [23, 24], E. I. Nochka [25], M. Ru [28, 29], B. Shiffman [30], B. Shabat [31], F. Sakai [32, 33], W. Stoll [37, 38], P. Vojta [36], H. Wu [43] and etc., and refer also to [1, 2, 3, 11, 12, 13] and references therein. To our knowledge, almost all the results in Nevanlinna theory are built just on Kähler manifolds which admit a complete Kähler metric of non-positive sectional curvature, while we seem to know little if the domain manifolds are not non-positively curved, because it is hard to estimate the local Green functions for bounded domains under other curvature conditions.
Recently, the author [14] obtained the first result for the value distribution of meromorphic mappings on a complete Kähler manifold with non-negative Ricci curvature by introducing a new method: global Green function method. Based on this method, a positive answer to the Picard’s problem can be given provided that is of maximal volume growth, i.e.,
@@
The original technique in [14] is to construct a family of relatively compact domains exhausting by using a minimal positive global Green function for Based on the asymptotic estimate of obtained by Colding-Minicozzi [8], we can give an optimal estimate for the local Green function for in terms of integral forms. As a result, the author established the Second Main Theorem of meromorphic mappings on Let us introduce the main contribution and key technique in [14]. Let be the minimal positive global Green function of for Colding-Minicozzi [8] (see [21] also) described the asymptotic behavior of : there exists a constant such that
where denotes the Riemannian distance function of from By means of this asymptotic estimate, the author defined a family of relatively compact domains:
@@
According to Lemma 3.5 in [14], for any there exists such that
holds for all satisfying Based on this estimate, the Calculus Lemma can be derived. When is conformally equivalent to since has non-negative Ricci curvature and maximal volume growth (see [21]). So, by employing the standard arguments, the author established the Second Main Theorem of meromorphic mappings from into a complex projective manifold (see [14], Theorem 3.10), which gives a Picard’s theorem:
@@
In this paper, we will consider a weaker volume growth condition:
@@
By Li-Yau’s estimate [22], there exist constants such that
@@ 1.2. Main results
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form Fix a reduced divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and see definition in Section 3.2 of this paper. Assume that is non-parabolic, by which we mean that carries a positive global Green function, i.e., it satisfies the volume growth condition
Let be the Ricci form of associated with the metric and be the canonical line bundle over The characteristic function of is defined by
The main result is the following Second Main Theorem.
@@
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
@@
Corollary 1.3 shows that each meromorphic function satisfying on must be a constant, if it omits 2 distinct values.
@@ 2.1. Volume Comparison Theorem
A space form is defined as a complete Riemannian manifold with constant sectional curvature. Given a simply-connected space form with constant sectional curvature of dimension Let denote the Riemannian volume of a geodesic ball with radius in Besides, let be a complete Riemannian manifold with Ricci curvature of dimension Fix a point Let denote the Riemannian volume of a geodesic ball centered at with radius
The classical volume comparison theorem by Bishop-Gromov (see [5]) can be stated as follows.
@@
When Calabi-Yau (see [35]) obtained a lower bound of
@@
Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
@@ 2.2. Estimates of Heat Kernels
Let be a complete Riemannian manifold of dimension with Laplace-Beltrami operator The heat kernel of is the minimal positive fundamental solution of the following heat equation
@@ 3.1. Construction of
Assume that is non-parabolic, which means that satisfies the volume growth condition
Hence, there exists a unique minimal positive global Green function of for i.e.,
@@
By as it is evident to see that is relatively compact for any and the sequence exhausts for According to Sard’s theorem, the boundary of is a submanifold of for almost every Set
@@ 3.2. Nevanlinna’s Functions
where is the Riemannian volume element of
Equip with the Kähler metric Let be the Hilbert-Schmidt norm of differential Since is Kählerian, then we have
@@ 3.3. First Main Theorem
To establish the First Main Theorem of we need Jensen-Dynkin formula (see, e.g., [12, 13]) as follows.
@@ 4. Gradient Estimates of Green Functions
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
@@ 5.1. Calculus Lemma
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to we see that for any there exists a subset of finite Lebesgue measure such that
@@ 5.2. Logarithmic Derivative Lemma
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
@@
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
@@ 6. Second Main Theorem and Defect Relation
Let be a non-compact complete Kähler manifold of complex dimension with non-negative Ricci curvature. Assume that is non-parabolic.
@@
where is the Ricci form of
Let be a divisor, where are prime divisors. The reduced form of is defined as
@@ 6.1. Second Main Theorem
in the sense of currents, where is the reduced form of Thus, it yields that
We establish a Second Main Theorem as follows.
@@
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that
@@ 6.2. Defect Relation
Endowing with Kähler metric Let be the Hilbert-Schmidt norm of differential with respect to metrics Let denote the scalar curvature of Take a local holomorphic coordinate the Ricci curvature tensor of can be written as the form
@@
Corollary 6.8 (Picard’s Theorem).
@@ 6.3. Picard’s Theorem
Let be a non-compact complete Kähler manifold with non-negative Ricci curvature. Let be distinct values in Let be a nonconstant meromorphic function on If is non-parabolic, then for any there exists a subset of finite Lebesgue measure such that